OpenStudy (anonymous):

Find a cubic function with the given zeros. √6 , -√6, -3

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what do we know about zeros of a function, what do they do to the function?

OpenStudy (anonymous):

they mark the intercepts?

OpenStudy (amistre64):

yeah, they do mark the x intercepts. but something more useful to is is that the make the function, equal out to zero.

OpenStudy (amistre64):

now, another thing we should know, is that anything times 0, is equal to 0 ... right?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

so, would you agree that if x=a, then x-a = 0 ?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

given any number of roots, we can then create a function: say the roots are a,b,c let f(x) = (x-a)(x-b)(x-c) will f(a)=0? f(b)=0? and f(c)=0?

OpenStudy (anonymous):

okay, so we have to find which one ?

OpenStudy (amistre64):

we simply need to expand (multiply thru, distribute) this into the cubic format.

OpenStudy (amistre64):

let f(x) = (x-a)(x-b)(x-c) for the stated zeros; a,b, and c how would we expand this into a cubic format?

OpenStudy (anonymous):

multiply it out

OpenStudy (amistre64):

can you show me how it looks when we multiply it out?

OpenStudy (amistre64):

that link is not working for my browser :/ (x-c)(xx-bx-ax+ba) xxx-bxx-axx+bax -cxx+bcx+cax-abc x^3 -(a+b+c)xx^2 +(ab+ac+bc)x -abc

OpenStudy (amistre64):

let a=sqrt(6), b=-sqrt(6), and c=-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

btw the link says x^3 + 3x^2 -6x -18

OpenStudy (amistre64):

a+b+c = -3, so --3 = 3 is fine ab+ac+bc = -6, thats good and -(--6.3) is -18

OpenStudy (anonymous):

so now we just combine them to get the function

OpenStudy (amistre64):

that IS the function ..

OpenStudy (amistre64):

f(x) = x^3 +3x^2-6x-18

OpenStudy (amistre64):

any scaled form of this also works ... so this is not a unique solution

OpenStudy (anonymous):

i see, thanks so much!

OpenStudy (amistre64):

goodluck